Extra Credit: Can you calculate and oxide weight percent for Corundum?
Answer #1

The molecular formula of corundum is Al2O3.

The percentage of O is calculated:

% O = 3 * AM O * 100 / MM Al2O3 = 3 * 16 * 100/102 = 47.1%

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.